OpenStudy (anonymous):

what slope equation would i use if all i know is the slope is (-9/2) and the point (8, -5) ?

OpenStudy (aroub):

-5-2/8+9 -7/17

OpenStudy (anonymous):

point slope formula

OpenStudy (anonymous):

y-y1=m(x-x1) m=slope y1=y value of the point given x1=x value of point given